OpenStudy (kropot72):

Question on series to follow:

OpenStudy (kropot72):

If \[a=b-\frac{b ^{2}}{2}+\frac{b ^{3}}{3}-\frac{b ^{4}}{4}+........\] express b in ascending powers of a.

OpenStudy (anonymous):

a = ln(b+1) b = -1 + e^a b = -1 + 1 + a + a^2/2! + a^3/3! + .......... b = a + a^2/2! + a^3/3! + .......... Derivation : For infinite G.P. 1 - b + b^2 - b^3 + .......... = 1/(1+b) For -1<=b<=1 integrate both sides Get ln(1+b) = b - b^2/2 + .....